# Pseudodifferential damping estimates and stability of relaxation shocks

### v3

@@ 8.1. Extensions and simplification
Our results yield immediately a number of simplifications and extensions for existing work. For example, In the 1-D large-amplitude relaxation case, they recover and somewhat simplify the damping result of [35], removing contortions associated with translating linear algebraic estimators to Friedrichs symmetrizer form. Likewise, in the viscous shock case [57, 58], global multi-D stability is known only for hyperbolic-parabolic systems of “Kawashima type” (roughly, simultaneously symmetrizable) at the endstates of the shock. Yet, high-frequency bounds (1.3) have been shown in the different context of bounded-time vanishing viscosity already in [16, 17] by Kreiss symmetrizer techniques under the much weaker, apparently sharp assumption [16, 17, (H5)] that the endstates be stable as constant solutions, a condition on the dispersion relation of the associated Fourier symbol. As simultaneous symmetrizability was used in [57, 58] only to establish a nonlinear damping estimate, we thus extend immediately the multi-D long-time stability results of [57, 58] to the larger class of systems considered in [16, 17], putting the two theories (long-time vs. small-viscosity) finally on par.
@@ 8.2. The small-amplitude case
Theorem 3.6 both simplifies and extends a number of small-amplitude results, replacing symmetrizability assumptions with the constant-coefficient high frequency damping condition (3.11), which is implied by but more general than the standard dissipativity conditions for the endstates of the wave, themselves implied by but more general than the symmetry-based Kawashima conditions of [23, 22, 56]. In particular, this extends the results of [25, 26] on stability of smooth multi-D relaxation profiles from symmetrizable to merely high-frequency dissipative systems. See [52] for a subsequent extension to stability of smooth multi-D shock profiles of second-order hyperbolic relaxation systems arising in relativistic gas dynamics.
@@ 8.3. The large-amplitude multi-D case
As regards our main motivating example, of large-amplitude multi-dimensional shocks in the Saint Venant model, the preliminary WKB analysis has been already done in [55, §2.4, p. 32] for the linear problem, yielding the high-frequency linearized spectral stability needed for our analysis, and verifying the assumptions made in Theorem 6.2. We note further, as pointed out in [55], that Airy points do appear for waves of all amplitudes, both smooth and discontinuous type, so that the analysis of Theorem 6.2 is indeed needed to obtain nonlinear damping. Combining these ingredients with the linearized resolvent bounds obtained in [25] for arbitrary-amplitude smooth multi-D profiles of general relaxation systems, one obtains the result that spectral stability in the sense of a standard Evans function condition [25, 55] implies linearized and nonlinear stability with standard decay rates matching those of the parabolic case [57, 58, 59]. Evans stability has been demonstrated for all such waves by nonrigorous numerics in [55]. It would be very interesting to convert this to rigorous computer-aided proof.

### v4

@@ 8.1. Extensions and simplification
Our results yield immediately a number of simplifications and extensions for existing work. For example, In the 1-D large-amplitude relaxation case, they recover and somewhat simplify the damping result of [38], removing contortions associated with translating linear algebraic estimators to Friedrichs symmetrizer form. Likewise, in the viscous shock case [62, 63], global multi-D stability is known only for hyperbolic-parabolic systems of “Kawashima type” (roughly, simultaneously symmetrizable) at the endstates of the shock. Yet, high-frequency bounds (1.3) have been shown in the different context of bounded-time vanishing viscosity already in [19, 20] by Kreiss symmetrizer techniques under the much weaker, apparently sharp assumption [19, 20, (H5)] that the endstates be stable as constant solutions, a condition on the dispersion relation of the associated Fourier symbol. As simultaneous symmetrizability was used in [62, 63] only to establish a nonlinear damping estimate, we thus extend immediately the multi-D long-time stability results of [62, 63] to the larger class of systems considered in [19, 20], putting the two theories (long-time vs. small-viscosity) finally on par.
@@ 8.2. The small-amplitude case
Theorem 3.6 both simplifies and extends a number of small-amplitude results, replacing symmetrizability assumptions with the constant-coefficient high frequency damping condition (3.11), which is implied by but more general than the standard dissipativity conditions for the endstates of the wave, themselves implied by but more general than the symmetry-based Kawashima conditions of [26, 25, 61]. In particular, this extends the results of [28, 29] on stability of smooth multi-D relaxation profiles from symmetrizable to merely high-frequency dissipative systems, assuming spectral stability. See [56] for a substantial further extension of these methods to the case of second-order hyperbolic relaxation systems arising in relativistic gas dynamics, yielding for the first time multi-D stability of smooth relativistic viscous shock profiles.
In the case considered in [60] of channel flow, one may verify spectral stability as well, to obtain a complete result independent of further assumptions. For [60], the Fourier frequencies $\xi$ now are selected from a discrete set $\xi_{j}=2\pi j/L$, $j\in{\mathbb{Z}}$, where $L$ is the width of the channel. For $|\xi_{j}|$ large, one obtains spectral stability from Theorem 3.6. For $|\xi_{j}|$ bounded and nonzero, one obtains spectral stability by continuity from the constant solution, which enjoys a spectral gap. For $\xi_{j}=0$, the spectral problem reduces to 1-D, hence stability follows by the 1-D result of [51]. This completes the proof of spectral stability. Nonlinear stability can then be shown by the essentially 1-D argument for flow in a channel developed in [57, 66], using exponential slaving of nonzero $\xi_{j}$ modes to the $\xi_{j}=0$ mode, together with the nonlinear damping estimate afforded by Theorem 3.6.
@@ 8.3. The large-amplitude multi-D case
As regards our main motivating example, of large-amplitude multi-dimensional shocks in the Saint Venant model, the preliminary WKB analysis has been already done in [60, §2.4, p. 32] for the linear problem, yielding the high-frequency linearized spectral stability needed for our analysis, and verifying the assumptions made in Theorem 6.2. We note further, as pointed out in [60], that Airy points do appear for waves of all amplitudes, both smooth and discontinuous type, so that the analysis of Theorem 6.2 is indeed needed to obtain nonlinear damping. Combining these ingredients with the linearized resolvent bounds obtained in [28] for arbitrary-amplitude smooth multi-D profiles of general relaxation systems, one obtains the result that spectral stability in the sense of a standard Evans function condition [28, 60] implies linearized and nonlinear stability with standard decay rates matching those of the parabolic case [62, 63, 64]. Evans stability has been demonstrated for all such waves by nonrigorous numerics in [60]. It would be very interesting to convert this to rigorous computer-aided proof.
@@ 9.1. Open problems
Finally, even for smooth multi-D relaxation profiles (on the whole space as opposed to a channel as discussed above), there remains the important open question of verifying multi-D spectral stability analytically in the small-amplitude limit, without resort to numerical approximation. As numerical Evans function computations become quite stiff in the small-amplitude limit, this is of practical as well as theoretical concern. See [51, 14, 15, 2] for models of such analyses in related situations.